OpenStudy (anonymous):

Factor the algebraic expression below in terms of a single trigonometric function. sin ^2x + sin x - 2

OpenStudy (campbell_st):

this is like factoring a quadratic expression \[\sin^2(x) + \sin(x) -2 = (\sin(x) +2)(\sin(x) -1)\]

OpenStudy (anonymous):

thats it ?

OpenStudy (campbell_st):

thats it its like factoring \[a^2 + a -2 = (a +2)(a -1)\]

OpenStudy (campbell_st):

as long at the expression is as I've written it.

OpenStudy (anonymous):

Thank you :)